OpenStudy (callisto):

How would you integrate this? \[\int (tanu+cotu)^2 du\]

OpenStudy (mimi_x3):

Well, expand it..

OpenStudy (callisto):

My attempt: \[\int (tanu+cotu)^2=\int (tan^2u+2+cot^2u) du\]\[=\int (sec^2u-1+2+csc^2u-1)= \int sec^2u +csc^2u du \]\[= tanu - cotu +C\] Any better method?

OpenStudy (callisto):

For the third step... I forgot to put du at the end:(

OpenStudy (mimi_x3):

lol, i did the same..was about to type that. xD

OpenStudy (callisto):

Oh...Thanks!!!!~

OpenStudy (mimi_x3):

Sorry I didnt help much..but im sure there's another method..

OpenStudy (callisto):

Nope!~ You did help!!! I always forget how to do this question :S